OpenStudy (anonymous):

find (fg)(2) and (f+g)(2) when f(x) = 1/x and g(x) = 4x + 9

OpenStudy (freckles):

(fg)(2)=f(2)*g(2) (f+g)(2)=f(2)+g(2)

OpenStudy (freckles):

so let's go ahead and find f(2) and g(2)

OpenStudy (freckles):

f(x)=1/x f(2)=1/2 find g(2)

OpenStudy (anonymous):

4(2) + 9 8 + 9 17

OpenStudy (freckles):

right no

OpenStudy (freckles):

so f(2)=1/2 and g(2)=17

OpenStudy (freckles):

so (fg)(2)=f(2)*g(2)=1/2 * 17

OpenStudy (freckles):

and (f+g)(2)=f(2)+g(2)=1/2+17

OpenStudy (freckles):

just apply the operations there and you are done

OpenStudy (anonymous):

Thank you!